OpenStudy (anonymous):

An outside circular ring has a circumference of 100 cm. What is the circumference of an inner ring which is 5 cm shorter in radius? Both circles have the same center.

OpenStudy (ganpat):

For Outer circle circumference = Perimeter = 2* Pi * r so, 100 = 2 * pi * R R = 15.91.. Inner circle radius is 5 cm less... so r= 15.91 - 5 = 10.91 and its circumference = 2 * pi * r = 2 * pi * 10.91 = 68.54 cm..

OpenStudy (anonymous):

so my answer is 68.54 ?

OpenStudy (ganpat):

yup.. understand the approach..

OpenStudy (anonymous):

yes i do thank you